OpenStudy (anonymous):

Please help!!!!!!!!!!!!! How can we prove that 1.9(and the nine is going on forever, and there is a line that is suppose to be on top of the 9) is equal to 2, without using rounding? Explain.

OpenStudy (anonymous):

Well,

OpenStudy (anonymous):

You can use division as proof...

OpenStudy (anonymous):

\[x=1.999...\\10x=19.999...\\9x=18\\x=2\]

OpenStudy (anonymous):

I don't really know how to show you though...

OpenStudy (anonymous):

I could show you with the logic of limits.

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

But before why did you start with 10? How did you get it?

OpenStudy (anonymous):

\[1.99999... = \lim_{^{x} \rightarrow \infty} 2-(0.1)^{x}\]

OpenStudy (anonymous):

@Rana12333 multiplying by 10 yields as follows...\[10x=1.999...\times10\\\ \ \ \ \ \ =19.999...\]Then, you may subtract from both sides.\[10x-x=19.999...-x\\9x=19.999...-1.999...\\\ \ \ \ =18\]Now, divide to isolate x.\[x=2\]

OpenStudy (anonymous):

Here, the expression (0.1)^x will approach 0

OpenStudy (anonymous):

Either odrin's or my explanation will work, but I actually like his better.

OpenStudy (anonymous):

But why did you chose the number 10?

OpenStudy (anonymous):

He chose 10 because that will shift the decimal point conveniently.

OpenStudy (anonymous):

I multiplied by 10 in order to do the proof.

OpenStudy (anonymous):

Can you explain it step by step so I can understand it better please

OpenStudy (anonymous):

I need 1.999999999999999999........, not 0.99999999999..........

OpenStudy (anonymous):

\[1.999... = 1+0.999...\]